simplify first and then solve the equation: x-(-6)=-4
Answer by checkley71(8403)

X-(-6)=-4 OR X+6=-4 OR X=-6-4 OR X=-10
PROOF -10-(-6)=-4 OR -10+6=-4 OR -4=-4